OpenStudy (anonymous):

given slope and point write slope intercept form of equation of the line through the given point with given slope through(-3,-5) slope= 1/2

jigglypuff314 (jigglypuff314):

slope intercept form is y = mx + b when slope = m so when given slope = 1/2 = m plug in so you get y = 1/2 x + b then you'll have to find b so temporarily plug in the given point (-3, -5) so -5 = (1/2)(-3) + b -5 = -3/2 + b b = -3.5 so you'll get y = 0.5x - 3.5

OpenStudy (anonymous):

or y = mx + b (-3,-5) x = -3 and y = -5 slope(m) = 1/2 now sub -5 = 1/2(-3) + b -5 = -3/2 + b -5 + 3/2 = b -10/2 + 3/2 = b -7/2 = b leaving you with the equation of : y = 1/2x - 7/2 Either way is correct, because -7/2 is the same as -3.5.

OpenStudy (anonymous):

thanks much both of you

jigglypuff314 (jigglypuff314):

glad we could help :)